zoe can rake her yard in 4 hours. stan can do the same job in only 3 hours. how long would it take zoe and stand to rake the yard if they worked together?
Answer by stanbon(75887)

zoe can rake her yard in 4 hours. stan can do the same job in only 3 hours. how long would it take zoe and stand to rake the yard if they worked together?
-------
zoe rate:: 1/4 job/hr
---
stan rate:: 1/3 job/hr
----
together rate:: 1/x job/hr
------
Equation:
rate + rate = together rate
1/4 + 1/3 = 1/x
3x + 4x = 4*3
7x = 12
x = 12/7 = 1.71 hrs = 1hr 0.71*60 min = 1hr 42sec (time to do the job together)
------
Cheers,
Stan H.